OpenStudy (anonymous):

(12/r+4) divide by (11/6r+24) simplify complex fraction

OpenStudy (anonymous):

\[\frac{12}{\frac{11 (r+4)}{6 r+24}}=\frac{12 (6 r+24)}{11 (r+4)}=\frac{72 (r+4)}{11 (r+4)}=\frac{72}{11} \]

OpenStudy (anonymous):

gracias finally figured how to give u medals jejeje :)